OpenStudy (anonymous):

Which of the following are quadratic functions? Check all that apply. A. y=2(x-3x^2+1) B. y=5+2x C. y=1/2x^2-13x+4 D. y=x/x-1

OpenStudy (anonymous):

do you know what a quadratic function is?

myininaya (myininaya):

is this a quadratic y=ax^2+bx+c? and can a be zero?

OpenStudy (anonymous):

no sir

OpenStudy (anonymous):

i know a cant be zero and that it has to be in the form y=ax^2+bx+c

myininaya (myininaya):

so which of the above can you write in the form y=ax^2+bx+c

OpenStudy (anonymous):

i think a and c is that correct?

myininaya (myininaya):

good!

OpenStudy (anonymous):

wooot! thanks

OpenStudy (anonymous):

are you sure C is a quadratic equation? look at the definition

myininaya (myininaya):

\[y=\frac{1}{2}x^2-13x+4\] is a quadratic

myininaya (myininaya):

is this not the function they have?

OpenStudy (anonymous):

oh. I thought the whole thing was on the denominator. but maybe not.

myininaya (myininaya):

maybe that is what the person meant but thats not how i read it